OpenStudy (anonymous):

Which equation could be used to find two consecutive odd integers whose sum is 32?

OpenStudy (anonymous):

1 3 5 7 difference between them is 2 odd integer = x second odd integer = x + 2

OpenStudy (anonymous):

x+x+2=32 pages would be 15 and 17 2x+2=32 2x=30